One Burger King Bacon Double Cheeseburger contains 33g of fat. Of this, 16g are saturated fats and the rest are monounsaturated and polyunsaturated fats. (A) If we assume that all of the saturated fats are stearic acid (18:0, aka octadecanoic acid, molecular weight 284.5 g/mol) molecules, how many ATP molecules and ATP equivalents are derived from the complete β oxidation of the saturated fats? Remember: For each mole of stearic acid undergoing β oxidation, 120 moles of ATP are generated, and Avogadro’s number = 6.02 x 1023 molecules/mol. (B) How many 400 amino acid-long proteins can be synthesized using the number of ATP and ATP equivalents calculated in part A? Include the ATPs required to charge each tRNA. (C) It has been estimated that there are 37.2 x 1012 cells in the human body, that there are 109 proteins per eukaryotic cell, and the average eukaryotic protein is 400 amino acids in length. Given your answer in B (and the fact that saturated fats account for ~50% of the fats in a Bacon Double Cheeseburger), how reasonable is it for one Bacon Double Cheeseburger to fulfill your dietary needs for protein synthesis?
Answer #1

1 mole of Stearic acid is equivalent to 284.5g of it.

A) Hence, 16g = 0.056 moles. One mole stearic acid gives 120 moles of ATP on oxidation.

So we have 0.056 * 120 = 6.72 moles of ATP = 6.72 * 6.02 * 1023 = 4.045 * 1024 numbers of ATP

B) for charging tRNA and loading a charged tRNA and translocating on ribosome requires 4 ATP. per Amino acid (AA)

So for 400 AA we need 1600 ATP. 4.045 * 1024 / 1600 = 2.528 * 1021 protein chains can be formed.

C) Total number of proteins in human body = 37.2 * 1012 * 109 = 4.0548 * 1015

Total ATP needed = 4.0548 *1015 * 1600 = 6.487 * 1018

Order difference of 106 is there. Hence Bacon Double Cheeseburger will fulfill the dietary needs of protein synthesis.